OpenStudy (biohazard9064):

Algebra Help. Solve -2(2x + 5) - 3 = -3(x - 1).

OpenStudy (lynfran):

-2(2x+5)-3=-3(x-1) -4x-10-3=-3x+3 -4x-13=-3x+3 -4x+3x=3+13 -x=16 x=-16

OpenStudy (biohazard9064):

Thanks